OpenStudy (anonymous):

an 8 ohm reSiStance, 16 ohm inductive reactance and an unknown capacitor are connected in SerieS acroSs a 100 Volt, 60 Hz Supply. the current drawn by the circuit iS found to be 12.5 Ampere. determine the value of the capacitance, the voltage acroSS the capacitor and p.f of the circuit

OpenStudy (anonymous):

you still here

OpenStudy (anonymous):

R=8 ohm wl=16 c=? V=100 V w=60 * 2pi \[\frac{V}{\sqrt{R^2+\left(\text{wl}-\frac{1}{\text{wc}}\right)}}=i\]

OpenStudy (anonymous):

\[\frac{100}{\sqrt{8^2+\left((16)-\frac{1}{(60*2\text{Pi})c}\right)^2}}\text{==}12.5,\]

OpenStudy (anonymous):

\[c -> 0.000165786\]